h + w + p
b = -------------, for w (a baseball formula)
a + w + p + f

I tried several methods but still can't solve it.

Answer by stanbon(75887)

h + w + p
b = -------------, for w (a baseball formula)
a + w + p + f
-------------------
b(a+w+p+f) = h+w+p
-------
ba + bw + bp + bf = h + w + p
-------
Get the "w" terms together::
w - bw = ba + bp + bf - h - p
-------
Factor out the "w":
w(1-b) = ba + bp +bf - h - p
-----
w = [ba + bp + bf - h - p]/[1-b)
------------------------
Cheers,
Stan H.
-----------------